# DriveMLM: Aligning Multi-Modal Large Language Models with Behavioral Planning States for Autonomous Driving

### v2

@@ 3.2 Behavioral Planning States Alignment
Transforming the linguistic choices of Large Language Models (LLMs) into actionable control signals is crucial for vehicle control. To achieve this, we align the LLM’s outputs with the decision stages of the behavioral planning module in the popular Apollo system. Following common practice [3], we divide the decision-making process into two categories: speed decision and path decision. Specifically, the speed decision states contain [KEEP, ACCELERATE, DECELERATE, STOP], while the path decision states include [FOLLOW, LEFT_CHANGE, RIGHT_CHANGE, LEFT_BORROW, RIGHT_BORROW].
To enable a language model to make precise predictions among these states, we established a comprehensive link between linguistic descriptions and decision states, as illustrated in System Massage of Table 1. This correlation is used as a part of the system message and is integrated into the MLLM planner. As a result, once the LLM describes certain situations, the prediction will converge into a clear decision within the decision space. At each time, one-speed decision and one path decision are mutually inferred and sent to the motion planning framework. More detailed definitions of decision states can be found in the supplementary material.
System Message: You are a driving assistant to drive the car. You need to follow the navigation command and traffic rules. The traffic rule is … Path decisions include [FOLLOW, …]. Path decision definitions: ‘FOLLOW’ means …, Speed decisions include [KEEP, …]. Speed decision definitions: ‘KEEP’ means …, Given navigation command and driving scene obtained from camera or LiDAR, You should choose a path decision and a speed decision from the predefined options and give the explanation of your decision. Q1(Human): Caption instruction (e.g. Describe the current driving environment.) A1(DriveMLM): Caption response (e.g. It is currently daytime. A red car is driving away in front of ego.) Q2(Human): Navigation instruction (e.g. The navigation command is turned right. Please choose a path decision and a speed decision.) A2(DriveMLM): Speed and path decision (e.g. RIGHT_CHANGE, KEEP) Q3(Human): Explanation instruction (e.g. Please explain why to choose these decisions.) A3(DriveMLM): Explanation response (e.g. Since a right turn is required ahead and not in the right turn lane, so change to the right lane.) Q4(Human): Instruction (e.g. I’m in a hurry. Can you overtake the front car?) A4(DriveMLM): Speed and path decision (e.g. LEFT_CHANGE, ACCELERATE) Q5(Human): Explanation instruction (e.g. Please explain why to choose these decisions.) A5(DriveMLM): Explanation response (e.g. Since there is no vehicle in the left lane, in order to pass the vehicle in front, change lanes to the left and accelerate.)
Table 1: Examples of system message and interaction between user and DriveMLM system. The system message includes the description of the driving task, the traffic rules, and the definition of decision states. Given driving scenes such as images and user prompts, the driving system can infer the image caption, path, and speed decision, and additional explanation. Complete system messages and prompts are provided in the supplementary.
@@ 3.3 MLLM Planner
Multi-Modal Tokenizer. This tokenizer is engineered to handle various forms of input efficiently:
(1) For temporal multi-view images: We use a temporal QFormer to process multi-view images from timestamp $-T$ to 0 (current timestamp). First, it takes each view $I_{i}^{-T}$ at timestamp $-T$ and feeds it to ViT-g and QFormer with $N_{Q}$ random initialized queries of $D$ dimension. This produces image token embedding $X_{I_{i}^{-T}}\!\in\!\mathbb{R}^{N_{Q}\times D}$. Then, using the image token embedding $X_{I_{i}}^{-T}$ as queries of QFormer, we get the image token embedding of the next timestamp $X_{I_{i}^{-T+1}}$ by conducting the first step again. We repeat the two steps until we get the image token embedding of the current timestamp $X_{{I_{i}}^{0}}$, which gathers all temporal information from $-T$ to 0. This approach avoids the linear increase in resources required to process time series data as the length of time increases.
(2) For LiDAR data, we first send the point clouds as the input of the Sparse Pyramid Transformer (SPT) backbone [74] to extract the LiDAR features. Then we employ Qformer with $M$ random initialized queries of $D$ dimension to get the point cloud token embedding $X_{L}\!\in\!\mathbb{R}^{N_{Q}\times D}$. We concatenate it with the image token embedding.
@@
MLLM Decoder. The decoder is the core that translates the tokenized inputs into decision states and decision explanations. To this end, we design a system message template for LLM-based AD, which is shown in Table 1. We see that the system messages contain a description of the AD tasks, traffic rules, the definition of decision states, and placeholders indicating where each modality’s information is incorporated. This approach ensures that inputs from various modalities and sources are seamlessly integrated.
The output is formatted to provide decision states (see the Q2 of Table 1) and an explanation of the decisions (see the Q3 of Table 1), offering transparency and clarity in the decision-making process. Regarding the supervision methods, our framework uses cross-entropy loss with the next token prediction, following common practices. In this way, the MLLM planner can perform detailed understanding and processing of data from different sensors and sources, and transform it into appropriate decisions and explanations.
@@ 4.1 Data Analysis
We have collected $280$ hours of driving data for training. These data consist of 50k routes, collected in 30 driving scenarios with different weather and lighting conditions across 8 maps (Town01, Town02, Town03, Town04, Town06, Town07, Town10HD, Town12) in CARLA. On average, each scenario has about 200 trigger points on each map to be randomly triggered. Each scenario is either a common or rare safety-critical situation in driving. Details of these scenarios are in the supplementary. For each frame, we collect images from 4 cameras on the front, rear, left, and right, and also the point clouds from a LiDAR sensor added in the center of the ego vehicle. All data we collected have corresponding explanations and accurate decisions that successfully drive through scenarios.
Table 2 presents the comparison with the previous datasets designed for driving understanding with natural language. Our data has two unique features. The first is the alignment of behavioral planning states. This enables us to transform the MLLM planner’s output to control signal so that our framework can control vehicles in closed-loop driving. The second is human interaction annotation. It is characterized by natural language instructions given by humans alongside the responding decisions and explanations. The objective is to improve the ability to understand human instructions and respond accordingly.
@@
NuPrompt [70] $\checkmark$
NuScenes-QA [52] $\checkmark$ $\checkmark$
Rank2Tell [55] $\checkmark$ $\checkmark$
BDD-X [72] $\checkmark$ $\checkmark$
DRAMA [44] $\checkmark$ $\checkmark$
DriveLM [16] $\checkmark$ $\checkmark$ $\checkmark$
Ours
@@
4.2 Implementation Details
Our MLLM model is built from LLaMA [23]. Specifically, we use ViT-g/14 from EVA-CLIP [20] as the visual encoder and LLaMA-7B [64] as the LLM. The querying transformer with $N_{Q}$ queries is applied to extract image tokens from ViT, where we set $N_{Q}=32$. For the LiDAR encoder, we use the GD-MAE [74] model finetuned on ONCE [42]. Based on the pre-trained husky model, we train MLLM with instruction following data. We employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a cosine learning rate decay with learning rate $5e^{-5}$. The training epoch is 2, and the batch size is 256. We train QFormer and LLM to ensure the instruction following ability of LLM so that we can obtain a predefined format of path decision and speed decision. The resolution of image input to MLLM is set as $448\times 448$.
[…]
follow change borrow keep accelerate decelerate stop
LLaVA 1.5 [37] LLM 22.92 0.73 0.00 0.00 0.75 0.00 0.02 0.00 10.00 18.03 23.00
InstructBLIP [17] LLM 17.92 0.00 0.30 0.08 0.23 0.00 0.28 0.00 9.81 18.61 22.95
Apollo [3] FSM 18.53 0.76 0.40 0.04 0.54 0.05 0.19 0.37 - - -
@@ 4.2 Implementation Details
For evaluating closed-loop driving performance, we use the widely used Town05Long benchmark, which follows previous work [15, 57]. It is worth noting that Town05 is not in our training data. We use Driving Score (DS), Route Completion (RC), and Infraction Score (IS) [18] as the metrics. RC computes the average percentage of routes completed by an agent. IS measures the infraction penalty between 00 and $1$, including collision and violation of traffic rules. Note that IS is only calculated on the completed part of a route. DS is the core metric among the three, which is the product of both RC and IS. We also evaluate driving performance using Miles Per Intervention (MPI), which is a widely used metric in industry. It is computed as the total distance traveled over the total times of human takeovers. If the ego-car violates traffic rules or has a collision, it will be taken over and continue self-driving in a safe location until it reaches its destination. Unlike DS, which terminates the route under certain conditions, MPI requires the ego-car to complete the entire route.
@@
For the open-loop evaluation, we collect 10 routes of each scenario in Town05 obtained and annotated by human drivers as the test set. To evaluate decision prediction, we compute the accuracy of predicted decision pairs and the F1 score of each type of decision. For the explanation prediction task, we use the commonly used metrics in the NLP community, including BLEU-4 [49], CIDEr [65] and METEOR [4]. We compare our method with the popular Apollo, which is based on a finite state machine (FSM) and two MLLM models - LLaVA1.5 [37] and InstructBLIP [17]. These two MLLM models used for comparison were not fine-tuned but instead provided with several examples of input/decision pairs for few-shot adaptation.
@@ 4.3 Evaluation in Closed-Loop Driving
Roach [81] DD 43.6 80.4 0.54 -
Interfuser [57] DD 68.3 95.0 0.72 0.70
ThinkTwice [27] DD 70.9 95.5 0.75 0.40
Apollo [3] FSM 71.4 92.2 0.80 0.76
@@
Table 4: Results of closed-loop driving on CARLA Town05 Long. FSM denotes a Finite State Machine. DD denotes Data Driven. DS denotes Driving Score. RC denotes Route Completion. IS denotes Infraction Score. MPI denotes Miles Per Intervention. DriveMLM has a higher driving score and route completion rate and is also close to Apollo’s infraction penalty, indicating that DriveMLM can make better decisions while following the traffic rules. Meanwhile, DriveMLM also shows advantages in MPI, representing fewer human takeovers at the same mileage.
We evaluate closed-loop driving in CARLA, the most widely used and realistic simulation benchmark publicly available. State-of-the-art methods [81, 57, 27] that are capable of performing closed-loop driving in CARLA are included for performance comparison. The open-sourced Apollo [3] is also evaluated in CARLA as a baseline. No other LLM-based methods have shown the readiness to be deployed and evaluated besides ours. All methods are evaluated on Town05 long benchmarks [15].
Table 4 presents the Driving Score, Route Completion, and Infraction Score. Note that despite being a rule-based method, Apollo achieves almost on-par performance with recent end-to-end methods. DriveMLM surpasses all other methods on Driving Score by a large margin. This suggests that DriveMLM is better for handling state-transitions to safely drive through hard cases. The last column in Table 4 presents the results of MPI evaluation. This metric shows a more holistic driving performance because an agent is required to finish all routes. In other words, all situations along all routes are encountered by the tested agents. Thinktwice achieves better DS but lower MPI than Interfuser due to frequently crossing the stop line. However, CARLA imposes minimal penalties for this behavior. By contrast, MPI takes each violation of traffic rules as one take-over. DriveMLM also achieves the highest MPI among all other methods, suggesting its ability to avoid more situations for a safer driving experience.
4.4 Evaluation of Driving Knowledge
We adopt open-loop evaluation to evaluate the driving knowledge, which includes the decision prediction and the explanation prediction task. Table 3 presents the accuracy of predicted decision pairs, F1-score of each type of decision for the decision prediction, and BLEU-4 [49], CIDEr [65] and METEOR [4] for the predicted explanation. For Apollo, manually collected scenarios on Town05 are replayed as input to models in Table 3. The corresponding model states and outputs at every timestamp of replay are saved as predictions for metric calculation. For other methods, we give them the corresponding images as input and the proper prompts. By comparing model prediction with our manually collected ground truth, accuracy reveals decision correctness and similarity to human behavior, and the F1-score demonstrates the decision-making capability across each individual type of path and speed decision. DriveMLM achieves the highest accuracy overall, surpassing LLaVA with an accuracy of 40.97%. Compared to the Apollo baseline, the higher F1-score of DriveMLM suggests that it is much more effective in overtaking the rule-based state machine for solving various road situations. LLaVA [37], InstructBLIP [17], and our proposed DriveMLM can output explanations of decisions in the form of question and answer. In terms of BLEU-4, CIDEr, and METEOR, DriveMLM can achieve the highest performance, indicating that DriveMLM can give the most reasonable explanation of the decision.

### v3

@@ 3.2 Behavioral Planning States Alignment
Transforming the linguistic choices of Large Language Models (LLMs) into actionable control signals is crucial for vehicle control. To achieve this, we align the LLM’s outputs with the decision stages of the behavioral planning module in the CARLA AutoPolot System.
We divide the decision-making process into two categories: speed and path decisions. Specifically, the speed decision states contain [KEEP, ACCELERATE, DECELERATE, STOP], while the path decision states include [FOLLOW, LEFT_CHANGE, RIGHT_CHANGE, LEFT_BORROW, RIGHT_BORROW].
To enable a language model to make precise predictions among these states, we established a comprehensive link between linguistic descriptions and decision states, as illustrated in System Massage of Fig. 4. This correlation is used as a part of the system message and is integrated into the MLLM planner. As a result, once the LLM describes certain situations, the prediction will converge into a clear decision within the decision space. At each time, one-speed decision and one path decision are mutually inferred and sent to the motion planning framework.
Figure 4: Examples of system message and interaction between user and DriveMLM system. The system message includes the description of the driving task, the traffic rules, and the definition of decision states. Given driving scenes such as images and user prompts, the driving system can infer the image caption, path, and speed decision, and additional explanation.
@@ 3.3 MLLM Planner
Multi-Modal Tokenizer. This tokenizer is engineered to handle three different forms of input efficiently.
(1) For temporal multi-view images: We use a temporal cross-attention model to process multi-view images from timestamp T-T to 0 (current timestamp). First, images from various perspectives and times are processed by the CLIP vision encoder to obtain image features. Next, these image features undergo a nonlinear transformation through a projector to be mapped to the same feature dimension as the text tokens. Finally, through the cross-attention layer, the current time’s image is used as the query, while the historical images are first flattened into tokens, and used as key and value inputs to the multi-head cross-attention layer to obtain the fused features.
(2) For LiDAR data, it is necessary to align LiDAR features with text features into the same feature space, which enables the model to understand and process data from different modalities in a unified manner. However, text descriptions of scenarios are difficult to obtain due to the limitation of related datasets, we used the image domain as an intermediary for connecting text and point cloud. An image-lidar CLIP model is applied here, where the image encoder is the image encoder of ViT-L/14 radford2021_clip , and we freeze all its parameters. For the LiDAR encoder, we utilized the single-stride sparse transformer (SST)fan2022embracing , which is randomly initialized. During the training process, an image is sent to the frozen image encoder to produce the target feature $X_{I}$, while the LiDAR feature is extracted by the SST, with the input of corresponding LiDAR data $X_{L}$. The LiDAR encoder is trained to maximize the similarity between $X_{L}$ and $X_{I}$ with a cosine similarity loss, for mapping the LiDAR features into image feature space. After obtaining the LiDAR feature $X_{L}$, we project it linearly to a LiDAR token $H_{L}$, which has the same feature dimensionality as the language tokens. We applied data pre-processing during training. Every LiDAR data point corresponds to several views of images, and we regard each of them as a corresponding pair. The point cloud is also transformed into the camera coordinate system, and we project the 3D points into 2D image space and drop all the points outside the boundary of images, which are invisible. During test, we doesn’t drop any points because there are multiple views for a timestep.
@@
MLLM Decoder. The decoder is the core that translates the tokenized inputs into decision states and decision explanations. To this end, we design a system message template for LLM-based AD, which is shown in Fig. 4. We see that the system messages contain a description of the AD tasks, traffic rules, the definition of decision states, and placeholders indicating where each modality’s information is incorporated. This approach ensures that inputs from various modalities and sources are seamlessly integrated.
The output is formatted to provide decision states (see the Q2 of Fig. 4) and an explanation of the decisions (see the Q3 of Fig. 4), offering transparency and clarity in the decision-making process. Regarding the supervision methods, our framework uses cross-entropy loss with the next token prediction, following common practices. In this way, the MLLM planner can perform detailed understanding and processing of data from different sensors and sources, and transform it into appropriate decisions and explanations.
@@ 4.1 Data Analysis
We have collected $280$ hours of driving data for training. These data consist of 50k routes, collected in 30 driving scenarios with different weather and lighting conditions across 8 maps (Town01, Town02, Town03, Town04, Town06, Town07, Town10HD, Town12) in CARLA. On average, each scenario has about 200 trigger points on each map to be randomly triggered. Each scenario is either a common or rare safety-critical situation in driving. For each frame, we collect images from 4 cameras on the front, rear, left, and right, and also the point clouds from a LiDAR sensor added in the center of the ego vehicle. All data has corresponding explanations and accurate decisions that successfully drive through scenarios.
Tab. 1 presents the comparison with the previous datasets designed for driving understanding with natural language. Our data has two unique features. The first is the alignment of behavioral planning states. This enables us to transform the MLLM planner’s output to control signal so that our framework can control vehicles in closed-loop driving. The second is human interaction annotation. It is characterized by natural language instructions given by humans alongside the responding decisions and explanations. The objective is to improve the ability to understand human instructions and respond accordingly.
@@
NuPrompt wu2023language $\checkmark$
NuScenes-QA qian2023nuscenes $\checkmark$ $\checkmark$
Rank2Tell sachdeva2023rank2tell $\checkmark$ $\checkmark$
BDD-X xu2023drivegpt4 $\checkmark$ $\checkmark$
DRAMA movva2023large $\checkmark$ $\checkmark$
DriveLM drivelm2023 $\checkmark$ $\checkmark$ $\checkmark$
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Our training data contains 30 common or rare safety-critical scenarios, and Tab. 2 lists the names of all the scenarios and describes the source of the scenarios. Non-custom scenarios (marked as $\dagger$ and $\ddagger$) are usually set by loading preset trigger points, which makes it difficult to set them in other maps. Therefore, we have dynamized the scenarios to automatically find suitable trigger points on any map in preparation for the scenario setup. It is worth noting that all scenarios in Tab. 2 have been dynamized.
   ID    Scenario Name
   1    YieldBehindEmergencyVehicles ($\star$)
   2    OvertakingFromLeft ($\star$)
   3    OvertakingFromRight ($\star$)
   4    LeftBorrowPassObstacle ($\dagger$)
   5    LeftBorrowPassAccident ($\dagger$)
   6    LeftInvasionBorrowPassObstacle ($\dagger$)
   7    LeftInvasionBorrowPassAccident ($\dagger$)
   8    RightBorrowPassObstacle ($\dagger$)
   9    RightBorrowPassAccident ($\dagger$)
   10    RightInvasionBorrowPassObstacle ($\dagger$)
   11    RightInvasionBorrowPassAccident ($\dagger$)
   12    JunctionRightChange ($\star$)
   13    JunctionLeftChange ($\star$)
   14    JunctionStraight ($\star$)
   15    JunctionYieldPedestrian ($\star$)
   16    JunctionYieldPedestrianAfterTurn ($\dagger$)
   17    YieldJunctionSpecialisedVehicles ($\dagger$)
   18    LeftChangeInRoute ($\star$)
   19    RightChangeInRoute ($\star$)
   20    UnprotectedJunctionLeftTurn ($\dagger$)
   21    UnprotectedJunctionStraight ($\dagger$)
   22    UnprotectedJunctionRightTurn ($\dagger$)
   23    SignedJunctionLeftTurn ($\dagger$)
   24    SignedJunctionStraight ($\dagger$)
   25    SignedJunctionRightTurn ($\dagger$)
   26    PedestrianBlindSpotA ($\ddagger$)
   27    PedestrianBlindSpotB ($\ddagger$)
   28    VehicleBlindSpotA ($\ddagger$)
   29    VehicleBlindSpotB ($\ddagger$)
   30    FollowerChange ($\dagger$)
Table 2: Scenario list. $\star$ denotes that these scenarios are constructed by ourselves. $\dagger$ and denotes that these scenarios are from official Carla settings and ReasonNetshao2023reasonnet , respectively.
Method Type Acc. $(\%)$ $\uparrow$ Path (F1) $\uparrow$ Speed (F1) $\uparrow$ BLEU-4 $\uparrow$ CIDEr $\uparrow$ METEOR $\uparrow$
follow change borrow keep accel. decel. stop
LLaVA 1.5 liu2023improved LLM 22.92 0.73 0.00 0.00 0.75 0.00 0.02 0.00 10.00 18.03 23.00
InstructBLIP instructblip LLM 17.92 0.00 0.30 0.08 0.23 0.00 0.28 0.00 9.81 18.61 22.95
Apollo apollo_auto FSM 18.53 0.76 0.40 0.04 0.54 0.05 0.19 0.37 - - -
@@ 30    FollowerChange ($\dagger$)
4.2 Implementation Details
Our MLLM model is built from LLaMA gao2023llama . Specifically, we use ViT-g/14 from EVA-CLIP fang2023eva as the visual encoder and LLaMA-7B touvron2023llama as the LLM. The querying transformer with $N_{Q}$ queries is applied to extract image tokens from ViT, where we set $N_{Q}=32$. For the LiDAR encoder, we use the GD-MAE yang2023gd model finetuned on ONCE mao2021one . Based on the pre-trained husky model, we train MLLM with instruction following data. We employ the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a cosine learning rate decay with learning rate $5e^{-5}$. The training epoch is 2, and the batch size is 256. We train QFormer and LLM to ensure the instruction following ability of LLM so that we can obtain a predefined format of path decision and speed decision. The resolution of image input to MLLM is set as $448\times 448$.
@@ 4.2 Implementation Details
For evaluating closed-loop driving performance, we use the widely used Town05Long benchmark, which follows previous work chitta2022transfuser ; shao2023safety . It is worth noting that Town05 is not in our training data. We use Driving Score (DS), Route Completion (RC), and Infraction Score (IS) dosovitskiy2017carla as the metrics. RC computes the average percentage of routes completed by an agent. IS measures the infraction penalty between 0 and $1$, including collision and violation of traffic rules. Note that IS is only calculated on the completed part of a route. DS is the core metric among the three, which is the product of both RC and IS. We also evaluate driving performance using Miles Per Intervention (MPI), which is a widely used metric in industry.It is computed as the total distance traveled over the total times of human takeovers. If the ego-car violates traffic rules or has a collision, it will be taken over and continue self-driving in a safe location until it reaches its destination. Unlike DS, which terminates the route under certain conditions, MPI requires the ego-car to complete the entire route.
For the open-loop evaluation, we collect 10 routes of each scenario in Town05 obtained and annotated by human drivers as the test set. To evaluate decision prediction, we compute the accuracy of predicted decision pairs and the F1 score of each type of decision.
@@
For the explanation prediction task, we use the commonly used metrics in the NLP community, including BLEU-4 papineni2002bleu , CIDEr vedantam2015cider and METEOR banerjee2005meteor . We compare our method with the popular Apollo, which is based on a finite state machine (FSM) and two MLLM models - LLaVA1.5 liu2023improved and InstructBLIP instructblip . These two MLLM models used for comparison were not fine-tuned but instead provided with several examples of input/decision pairs for few-shot adaptation.
4.3 Evaluation of Driving Knowledge
We adopt open-loop evaluation to evaluate the driving knowledge, which includes the decision prediction and the explanation prediction task. Tab. 3 presents the accuracy of predicted decision pairs, F1-score of each type of decision for the decision prediction, and BLEU-4 papineni2002bleu , CIDEr vedantam2015cider and METEOR banerjee2005meteor for the predicted explanation. For Apollo, manually collected scenarios on Town05 are replayed as input to models in Tab. 3. The corresponding model states and outputs at every timestamp of replay are saved as predictions for metric calculation. For other methods, we give them the corresponding images as input and the proper prompts. By comparing model prediction with our manually collected ground truth, accuracy reveals decision correctness and similarity to human behavior, and the F1-score demonstrates the decision-making capability across each individual type of path and speed decision. DriveMLM achieves the highest accuracy overall, surpassing LLaVA with an accuracy of 40.97%. Compared to the Apollo baseline, the higher F1-score of DriveMLM suggests that it is much more effective in overtaking the rule-based state machine for solving various road situations. LLaVA liu2023improved , InstructBLIP instructblip , and our proposed DriveMLM can output explanations of decisions in the form of question and answer. In terms of BLEU-4, CIDEr, and METEOR, DriveMLM can achieve the highest performance, indicating that DriveMLM can give the most reasonable explanation of the decision.
@@ 4.4 Evaluation in Closed-Loop Driving
We evaluate closed-loop driving in CARLA, the most widely used and realistic simulation benchmark publicly available. State-of-the-art methods zhang2021end ; shao2023safety ; jia2023think that are capable of performing closed-loop driving in CARLA are included for performance comparison. The open-sourced Apollo apollo_auto is also evaluated in CARLA as a baseline. No other LLM-based methods have shown the readiness to be deployed and evaluated besides ours. All methods are evaluated on Town05 long benchmarks chitta2022transfuser .
Tab. 4 presents the Driving Score, Route Completion, and Infraction Score. Note that despite being a rule-based method, Apollo achieves almost on-par performance with recent end-to-end methods. DriveMLM surpasses all other methods on Driving Score by a large margin.This suggests that DriveMLM is better for handling state-transitions to safely drive through hard cases. The last column in Tab. 4 presents the results of MPI evaluation. This metric shows a more holistic driving performance because an agent is required to finish all routes. In other words, all situations along all routes are encountered by the tested agents. Thinktwice achieves better DS but lower MPI than Interfuser due to frequently crossing the stop line. However, CARLA imposes minimal penalties for this behavior. By contrast, MPI takes each violation of traffic rules as one take-over. DriveMLM also achieves the highest MPI among all other methods, suggesting its ability to avoid more situations for a safer driving experience.
@@
Roach zhang2021end $D$ 43.6 80.4 0.54 -
Interfuser shao2023safety $D$ 68.3 95.0 0.72 0.70
ThinkTwice jia2023think $D$ 70.9 95.5 0.75 0.40
Apollo apollo_auto FSM 71.4 92.2 0.80 0.76
@@
Table 4: Results of closed-loop driving on CARLA Town05 Long. FSM denotes a Finite State Machine. $D$ denotes Data Driven. DS denotes Driving Score. RC denotes Route Completion. IS denotes Infraction Score. MPI denotes Miles Per Intervention. DriveMLM has a higher driving score and route completion rate and is also close to Apollo’s infraction penalty, indicating that DriveMLM can make better decisions while following the traffic rules. Meanwhile, DriveMLM also shows advantages in MPI, representing fewer human takeovers at the same mileage.